# Non-Hermitian Linear Electro-Optic Effect Through Interactions of Free and Bound Charges

### v1

@@ III.1 Description of the model
We consider a hypothetical material, schematically represented in Fig. 1, that consists of free carriers and bound charges in interaction.
@@
$m_{b}\left(\partial_{tt}+\gamma\partial_{t}+\sum_{j=x,y,z}\omega_{0j}^{2}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}\right)\mathbf{P}=q^{2}n_{b}\left(\mathbf{E}+\mathbf{\overline{C}}_{21}\cdot\frac{qn_{0}}{m}\mathbf{p}\right)$ (12)
@@
In the above, $\mathbf{P}$ is the polarization vector associated with the bound charges, $\omega_{0j}$ are the resonance frequencies (transverse optical phonon frequencies) along each axis, $\gamma$ is the phonon collision frequency, $m_{b}$ is the effective mass of the bound charges and $n_{b}$ their density. The notations $\partial_{t}\equiv\frac{\partial}{\partial t}$ and $\partial_{tt}\equiv\frac{\partial^{2}}{\partial t^{2}}$ are used for brevity. The vectors $\hat{\mathbf{e}}_{x}$, $\hat{\mathbf{e}}_{y}$, $\hat{\mathbf{e}}_{z}$ are the unit vectors along the Cartesian axes and $\otimes$ represents the tensor product of two vectors. For simplicity, we neglect the effect of the magnetic component of the Lorentz force in Eq. (11), which is justified when the drift velocity of the free carriers satisfies $|\mathbf{v}_{0}|\ll c$. The quasi-momentum distribution $\mathbf{p}$ of the free charges is related to the drift velocity as in Eq. (2).
In our model, the free carriers and the bound charges are coupled via two generic tensors, $\mathbf{\overline{C}}_{12}$ and $\mathbf{\overline{C}}_{21}$, with units of $\mathrm{Vm/A}$. The physical origin of the coupling is the following: the current densities of bound charges $\partial_{t}\mathbf{P}$ and of free carriers $qn_{0}\mathbf{v}\approx\frac{qn_{0}}{m}\mathbf{p}$ induced by an external field $\mathbf{E}$ generate secondary local electric fields given by $\mathbf{\overline{C}}_{12}\cdot\partial_{t}\mathbf{P}$ and $\mathbf{\overline{C}}_{21}\cdot\frac{qn_{0}}{m}\mathbf{p}$, respectively. These secondary local fields give rise to electric forces that interact with the free and bound charges, and that are at the origin of the coupling. The role of the tensors $\mathbf{\overline{C}}_{12}$ and $\mathbf{\overline{C}}_{21}$ is analogous to the Green’s function in electromagnetism, which mathematically translates the current densities (source terms) into radiated electric fields Kong (1986). Physically, the coupling between the bound and free charges depends on the actual microscopic distribution of charges and fields. Thus, $\mathbf{\overline{C}}_{12}$ and $\mathbf{\overline{C}}_{21}$ may be functions of the different dynamic variables of the system.
@@
In the absence of coupling, $\mathbf{\overline{C}}_{12}=\mathbf{\overline{C}}_{21}=0$, and of an anomalous velocity term, $\mathbf{\overline{\zeta}}=0$, equations (11) and (12) reduce to the standard Drude and Lorentz models. For completeness, in Appendix B we present a slightly refined version of the proposed model that ensures that the conduction current satisfies the continuity equation. As discussed in Appendix B, the corrections arising from such refinement are usually negligible.
@@ III.2 Allowed form of nonlinear coupling
$\frac{\partial}{{\partial t}}\left[{\frac{{{m_{b}}}}{{2{n_{b}}{q^{2}}}}{\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}+\frac{m_{b}}{{2{n_{b}}{q^{2}}}}\sum\limits_{j}{\omega_{0j}^{2}{\bf{P}}\cdot{{{\bf{\hat{e}}}}_{j}}\otimes{{{\bf{\hat{e}}}}_{j}}}\cdot{\bf{P}}}\right]$ (14)
$+\gamma\frac{{{m_{b}}}}{{{n_{b}}{q^{2}}}}{\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}={\bf{E}}\cdot{\partial_{t}}{\bf{P}}+{\partial_{t}}{\bf{P}}\cdot{\overline{\bf{C}}_{21}}\cdot\frac{{q{n_{0}}}}{m}{\bf{p}}$
On the other hand, from the Poynting theorem we have that $\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\rm{EM}}}}=-{\bf{E}}\cdot{{\bf{j}}_{{\rm{tot}}}}$, where $W_{\rm{EM}}$ is defined as in Sec. II.1 and ${{\bf{j}}_{{\rm{tot}}}}$ is the total current coupled to Maxwell’s equations. It is now given by the sum of the conduction and polarization currents: ${{\bf{j}}_{{\rm{tot}}}}={n_{0}}q{\bf{v}}+{\partial_{t}}{\bf{P}}$. Using Eqs. (13)-(14), it is readily found that:
$\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\rm{tot}}}}=-Q+\frac{{q{n_{0}}}}{m}\left[{{\partial_{t}}{\bf{P}}\cdot{{\overline{\bf{C}}}_{21}}\cdot{\bf{p}}+{\bf{p}}\cdot{{\overline{\bf{C}}}_{12}}\cdot{\partial_{t}}{\bf{P}}}\right]$ (15)
@@
${W_{{\rm{tot}}}}$ $={W_{{\rm{EM}}}}+{\frac{{{n_{0}}}}{{2m}}{\bf{p}}\cdot{\bf{p}}}+\frac{{{m_{b}}}}{{2{n_{b}}{q^{2}}}}\left({\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}+\sum\limits_{j}{\omega_{0j}^{2}{\bf{P}}\cdot{{{\bf{\hat{e}}}}_{j}}\otimes{{{\bf{\hat{e}}}}_{j}}}\cdot{\bf{P}}\right),$ (16)
$Q$ $=\frac{{{n_{0}}\Gamma}}{m}{\bf{p}}\cdot{\bf{p}}+\gamma\frac{{{m_{b}}}}{{{n_{b}}{q^{2}}}}{\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}.$ (17)
The terms of $W_{\rm{tot}}$ represent the electromagnetic field energy density, the energy density of the free carriers, and the kinetic and potential energy density of the bound charges, respectively.
@@
where the superscript $T$ represents the transpose operator. When the above constraint is fulfilled, Eq. (15) simplifies to the standard law of energy conservation Pitaevskii and Lifshitz (2012) with a light and a matter part $\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\rm{tot}}}}=-Q$, where $Q>0$ represents the heat dissipated per unit of volume. The condition (18) will be assumed in the rest of the article.
@@ III.3 Microscopic reversibility
The vast majority of physical systems obey the principle of microscopic reversibility, i.e. they are time-reversal (TR) invariant at the microscopic level Sachs (1987); Asadchy et al. (2020a). Here, we impose that the coupling terms in Eqs. (11) and (12) are compatible with such a principle and remain invariant under a TR symmetry transformation. The terms $\mathbf{\overline{C}}_{21}\cdot\mathbf{p}$ and $-\mathbf{\overline{C}}_{21}^{T}\cdot\partial_{t}\mathbf{P}$ must transform as “forces”, and thus must be even under a TR symmetry transformation (similar to the other terms in the equations, with exception of the dissipative terms). Because $\mathbf{p}$ and $\partial_{t}\mathbf{P}$ are both odd under a TR symmetry transformation, it follows that $\mathbf{\overline{C}}_{21}$ must also be odd under a TR, implying that $\mathbf{\overline{C}}_{21}$ cannot be a constant matrix. Thus, to respect microscopic reversibility, $\mathbf{\overline{C}}_{21}$ has to depend only on the state variables that are odd under a TR, namely: $\mathbf{v}$, $\partial_{t}\mathbf{P}$, $\partial_{t}\mathbf{E}$ or $\mathbf{B}$. In particular, it follows that the coupling terms associated with the tensors $\mathbf{\overline{C}}_{12}\cdot\partial_{t}\mathbf{P}$ and $\mathbf{\overline{C}}_{21}\cdot\mathbf{p}$ represent nonlinear interactions between the bound and free charges. Such a nonlinearity is essential to generate the NHEO effect Lannebère et al. (2022). We will demonstrate later in Sec. IV that the linearization of these nonlinear terms at the bias point may lead to a nonreciprocal and non-Hermitian electromagnetic response, even in the absence of a Berry curvature dipole $\left(\mathbf{\overline{\zeta}}=0\right)$.
@@
where the dummy indices $i$, $j$ and $l$ run over $x,y,z$ and the coupling coefficients $a_{ijl}$ and $b_{ijl}$ are real-valued coefficients (54 in total) with units of $\mathrm{Vs/A}$ and $\mathrm{m^{3}V/A^{2}}$, respectively. Note that in the lossless limit ($\gamma\to 0^{+}$ and $\Gamma\to 0^{+}$) the full dynamical equations become TR-symmetric.
@@ III.4 Linearization
$m_{b}\left(\partial_{tt}+\gamma\partial_{t}+\sum_{j}\omega_{0j}^{2}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}\right)\mathbf{P}$ $=q^{2}n_{b}\left(\mathbf{E}+\sum_{ijl}\left(a_{ijl}\hat{\mathbf{e}}_{l}\cdot\mathbf{v}+b_{ijl}\hat{\mathbf{e}}_{l}\cdot\partial_{t}\mathbf{P}\right)\hat{\mathbf{e}}_{i}\otimes\hat{\mathbf{e}}_{j}\cdot qn_{0}\mathbf{v}\right)$ (20b)
@@
${\bf{P}}_{\omega}^{0}=\varepsilon_{0}\sum\limits_{j}{\frac{\omega_{b}^{2}}{{\omega_{0j}^{2}-\omega\left({\omega+i\gamma}\right)}}{{{\bf{\hat{e}}}}_{j}}\otimes{{{\bf{\hat{e}}}}_{j}}}\cdot{{\bf{E}}_{\omega}}$ (21b)
where $\omega_{b}^{2}=\frac{q^{2}n_{b}}{\varepsilon_{0}m_{b}}$. In particular, the response to a static bias (${\bf{E}}={{\bf{E}}_{0}}$) is given by:
@@
$\mathbf{P}_{0}$ $=\varepsilon_{0}\sum_{j}\frac{\omega_{b}^{2}}{\omega_{0j}^{2}}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}\cdot\mathbf{E}_{0}.$ (22b)
@@
$\hskip 85.35826pt\varepsilon_{0}\omega_{b}^{2}\left({\mathbf{E}_{\omega}}+qn_{0}\sum_{jlr}v_{0r}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{l}\cdot\left[\left(a_{jlr}+a_{jrl}\right){\mathbf{v}_{\omega}}-i\omega b_{jrl}{\mathbf{P}_{\omega}}\right]\right)$
@@
$\mathbf{\overline{\varepsilon}}=\mathbf{\overline{\varepsilon}}_{\rm{L}}+\mathbf{\overline{\chi}}_{\rm{f-b}}+\mathbf{\overline{\chi}}_{\rm{f-f}},$ (25)
where $\mathbf{\overline{\varepsilon}}_{\rm{L}}$ determines the material response without the bias,
$\mathbf{\overline{\varepsilon}}_{\rm{L}}$ $=\sum_{j}\varepsilon_{jj}^{0}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}$ (26a)
$\varepsilon_{jj}^{0}(\omega)$ $=1-\frac{\omega_{p}^{2}}{\omega\left(\omega+i\Gamma\right)}+\frac{\omega_{b}^{2}}{\omega_{0j}^{2}-\omega^{2}-i\omega\gamma}$ (26b)
and $\mathbf{\overline{\chi}}_{\rm{f-b}}$ and $\mathbf{\overline{\chi}}_{\rm{f-f}}$ represent the electro-optic response due to the interaction of bound and free electrons and due to the “anomalous velocity”, respectively:
$\mathbf{\overline{\chi}}_{\rm{f-b}}$ $=-i\sum_{jlr}\left[g_{lr}a_{ljr}-g_{jr}\left(a_{jlr}+a_{jrl}\right)-f_{jlr}b_{jrl}\right]\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{l}$ (27a)
$\mathbf{\overline{\chi}}_{\rm{f-f}}$ $=\frac{{{{\mathbf{\overline{\sigma}}}_{{\rm{EO}}}}}}{{-i\omega{\varepsilon_{0}}}}.$ (27b)
@@
$f_{jlr}(\omega)$ $=-\varepsilon_{0}qn_{0}v_{0r}\omega\frac{\omega_{b}^{2}}{\omega_{0j}^{2}-\omega\left(\omega+i\gamma\right)}\frac{\omega_{b}^{2}}{\omega_{0l}^{2}-\omega\left(\omega+i\gamma\right)},$ (28a)
$g_{jr}(\omega)$ $=\varepsilon_{0}v_{0r}\frac{\omega_{p}^{2}\omega_{b}^{2}}{\left(\omega+i\Gamma\right)\left[\omega_{0j}^{2}-\omega\left(\omega+i\gamma\right)\right]}.$ (28b)
The linearized permittivity (25) is a dispersive function that in the absence of coupling ($a_{ijk}=b_{ijk}=0$ for all indices and $\mathbf{\overline{\zeta}}=0$) reduces to a diagonal matrix. As seen, the electro-optic effect may induce non-diagonal terms in the permittivity tensor. Remarkably, as discussed in the next section, for low-symmetry materials such non-diagonal terms may result in a permittivity tensor lacking transpose symmetry and may give rise to nonreciprocity and optical gain. In the rest of the article, we focus our analysis in the free and bound electrons interactions, and thereby we will ignore possible anomalous velocity contributions $\left(\mathbf{\overline{\zeta}}=0\right)$.

### v2

@@ III.1 Description of the model
We consider a low-symmetry conductor, schematically represented in Fig. 1, that consists of free carriers and bound charges in interaction.
@@
$m_{\text{b}}\left(\partial_{tt}+\gamma\partial_{t}+\sum_{j=x,y,z}\omega_{0j}^{2}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}\right)\mathbf{P}=q^{2}n_{\text{b}}\left(\mathbf{E}+\mathbf{\overline{C}}_{21}\cdot\frac{qn_{0}}{m}\mathbf{p}\right)$ (12)
@@
In the above, $\mathbf{P}$ is the polarization vector associated with the bound charges, $\omega_{0j}$ are the resonance frequencies (transverse optical phonon frequencies) along each axis, $\gamma$ is the phonon collision frequency, $m_{\text{b}}$ is the effective mass of the bound charges and $n_{\text{b}}$ their density. The notations $\partial_{t}\equiv\frac{\partial}{\partial t}$ and $\partial_{tt}\equiv\frac{\partial^{2}}{\partial t^{2}}$ are used for brevity. The vectors $\hat{\mathbf{e}}_{x}$, $\hat{\mathbf{e}}_{y}$, $\hat{\mathbf{e}}_{z}$ are the unit vectors along the Cartesian axes and $\otimes$ represents the tensor product of two vectors. For simplicity, we neglect the effect of the magnetic component of the Lorentz force in Eq. (11), which is justified when the drift velocity of the free carriers satisfies $|\mathbf{v}_{0}|\ll c$. The quasi-momentum distribution $\mathbf{p}$ of the free charges is related to the drift velocity as in Eq. (2).
@@
In our model, the free carriers and the bound charges are coupled via two generic tensors, $\mathbf{\overline{C}}_{12}$ and $\mathbf{\overline{C}}_{21}$, with units of $\mathrm{Vm/A}$. In the absence of coupling, $\mathbf{\overline{C}}_{12}=\mathbf{\overline{C}}_{21}=0$, and of an anomalous velocity term, $\mathbf{\overline{\zeta}}=0$, equations (11) and (12) reduce to the standard Drude and Lorentz models. For completeness, in Appendix B we present a slightly refined version of the proposed model that ensures that the conduction current satisfies the continuity equation. As discussed in Appendix B, the corrections arising from such refinement are usually negligible.
The physical origin of the nonlinear coupling can be understood as follows: the current densities of bound charges $\partial_{t}\mathbf{P}$ and of free carriers $qn_{0}\mathbf{v}\approx\frac{qn_{0}}{m}\mathbf{p}$ induced by an external field $\mathbf{E}$ generate secondary local electric fields given by $\mathbf{\overline{C}}_{12}\cdot\partial_{t}\mathbf{P}$ and $\mathbf{\overline{C}}_{21}\cdot\frac{qn_{0}}{m}\mathbf{p}$, respectively. These secondary local fields give rise to forces (e.g., through the nonlinearity of the Lorentz force or nonclassical effects) that interact with the free and bound charges. Physically, the coupling between the bound and free charges depends on the actual microscopic distribution of charges and fields. Thus, $\mathbf{\overline{C}}_{12}$ and $\mathbf{\overline{C}}_{21}$ may be functions of the different dynamic variables of the system.
It is relevant to note that previous works have considered nonlinear effects in the material response of metals Scalora et al. (2010); Ciracì et al. (2013), including interactions between bound and free electrons. However, these typically involve couplings mediated by the Lorentz force, as in microscopic field theory. In contrast, our model introduces a more general class of nonlinear couplings, permissible within the framework of macroscopic electrodynamics, where the fields are effective and may include additional interaction terms arising from the peculiar physics of Bloch electrons.
@@ III.2 Allowed form of nonlinear coupling
$\frac{\partial}{{\partial t}}\left[{\frac{{{m_{\text{b}}}}}{{2{n_{\text{b}}}{q^{2}}}}{\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}+\frac{m_{\text{b}}}{{2{n_{\text{b}}}{q^{2}}}}\sum\limits_{j}{\omega_{0j}^{2}{\bf{P}}\cdot{{{\bf{\hat{e}}}}_{j}}\otimes{{{\bf{\hat{e}}}}_{j}}}\cdot{\bf{P}}}\right]$ (14)
$+\gamma\frac{{{m_{\text{b}}}}}{{{n_{\text{b}}}{q^{2}}}}{\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}={\bf{E}}\cdot{\partial_{t}}{\bf{P}}+{\partial_{t}}{\bf{P}}\cdot{\overline{\bf{C}}_{21}}\cdot\frac{{q{n_{0}}}}{m}{\bf{p}}$
On the other hand, from the Poynting theorem we have that $\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\text{EM}}}}=-{\bf{E}}\cdot{{\bf{j}}_{{\text{tot}}}}$, where $W_{\text{EM}}$ is defined as in Sec. II.1 and ${\bf{j}}_{{\text{tot}}}$} is the total current coupled to Maxwell’s equations. It is now given by the sum of the conduction and polarization currents: ${{\bf{j}}_{{\text{tot}}}}={n_{0}}q{\bf{v}}+{\partial_{t}}{\bf{P}}$. Using Eqs. (13)-(14), it is readily found that:
$\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\text{tot}}}}=-Q+\frac{{q{n_{0}}}}{m}\left[{{\partial_{t}}{\bf{P}}\cdot{{\overline{\bf{C}}}_{21}}\cdot{\bf{p}}+{\bf{p}}\cdot{{\overline{\bf{C}}}_{12}}\cdot{\partial_{t}}{\bf{P}}}\right]$ (15)
@@
${W_{{\text{tot}}}}$ $={W_{{\text{EM}}}}+{\frac{{{n_{0}}}}{{2m}}{\bf{p}}\cdot{\bf{p}}}+\frac{{{m_{\text{b}}}}}{{2{n_{\text{b}}}{q^{2}}}}\left({\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}+\sum\limits_{j}{\omega_{0j}^{2}{\bf{P}}\cdot{{{\bf{\hat{e}}}}_{j}}\otimes{{{\bf{\hat{e}}}}_{j}}}\cdot{\bf{P}}\right),$ (16)
$Q$ $=\frac{{{n_{0}}\Gamma}}{m}{\bf{p}}\cdot{\bf{p}}+\gamma\frac{{{m_{\text{b}}}}}{{{n_{\text{b}}}{q^{2}}}}{\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}.$ (17)
The terms of $W_{\text{tot}}$ represent the electromagnetic field energy density, the energy density of the free carriers, and the kinetic and potential energy density of the bound charges, respectively.
@@
where the superscript $T$ represents the transpose operator. When the above constraint is fulfilled, Eq. (15) simplifies to the standard law of energy conservation Pitaevskii and Lifshitz (2012) with a light and a matter part $\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\text{tot}}}}=-Q$, where $Q>0$ represents the heat dissipated per unit of volume. The condition (18) will be assumed in the rest of the article.
@@ III.3 Microscopic reversibility
The vast majority of physical systems obey the principle of microscopic reversibility, i.e. they are time-reversal (TR) invariant at the microscopic level Sachs (1987); Asadchy et al. (2020a). Here, we impose that the coupling terms in Eqs. (11) and (12) are compatible with such a principle and remain invariant under a TR symmetry transformation. The terms $\mathbf{\overline{C}}_{21}\cdot\mathbf{p}$ and $\mathbf{\overline{C}}_{21}^{T}\cdot\partial_{t}\mathbf{P}$ must transform as “forces”, and thus must be even under a TR symmetry transformation (similar to the other terms in the equations, with exception of the dissipative terms). Because $\mathbf{p}$ and $\partial_{t}\mathbf{P}$ are both odd under a TR symmetry transformation, it follows that $\mathbf{\overline{C}}_{21}$ must also be odd under a TR, implying that $\mathbf{\overline{C}}_{21}$ cannot be a constant matrix. Thus, to respect microscopic reversibility, $\mathbf{\overline{C}}_{21}$ has to depend only on the state variables that are odd under a TR, namely: $\mathbf{v}$, $\partial_{t}\mathbf{P}$, $\partial_{t}\mathbf{E}$ or $\mathbf{B}$. In particular, it follows that the coupling terms associated with the tensors $\mathbf{\overline{C}}_{12}\cdot\partial_{t}\mathbf{P}$ and $\mathbf{\overline{C}}_{21}\cdot\mathbf{p}$ represent nonlinear interactions between the bound and free charges. We will demonstrate later in Sec. IV that the linearization of these nonlinear terms at the bias point may lead to a nonreciprocal and non-Hermitian electromagnetic response, even in the absence of a Berry curvature dipole $\left(\mathbf{\overline{\zeta}}=0\right)$.
@@
where the dummy indices $i$, $j$ and $l$ run over $x,y,z$ and the coupling coefficients $a_{ijl}$ and $b_{ijl}$ are real-valued coefficients (54 in total) with units of $\mathrm{Vs/A}$ and $\mathrm{m^{3}V/A^{2}}$, respectively. Note that in the lossless limit ($\gamma\to 0^{+}$ and $\Gamma\to 0^{+}$) the full dynamical equations become TR-symmetric. The parameters in this phenomenological model are understood as effective quantities, analogous to the effective mass in band theory.
@@ III.4 Linearization
$m_{\text{b}}\left(\partial_{tt}+\gamma\partial_{t}+\sum_{j}\omega_{0j}^{2}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}\right)\mathbf{P}$ $=q^{2}n_{\text{b}}\left(\mathbf{E}+\sum_{ijl}\left(a_{ijl}\hat{\mathbf{e}}_{l}\cdot\mathbf{v}+b_{ijl}\hat{\mathbf{e}}_{l}\cdot\partial_{t}\mathbf{P}\right)\hat{\mathbf{e}}_{i}\otimes\hat{\mathbf{e}}_{j}\cdot qn_{0}\mathbf{v}\right)$ (20b)
@@
${\bf{P}}_{\omega}^{0}=\varepsilon_{0}\sum\limits_{j}{\frac{\omega_{\text{b}}^{2}}{{\omega_{0j}^{2}-\omega\left({\omega+i\gamma}\right)}}{{{\bf{\hat{e}}}}_{j}}\otimes{{{\bf{\hat{e}}}}_{j}}}\cdot{{\bf{E}}_{\omega}}$ (21b)
where $\omega_{\text{b}}^{2}=\frac{q^{2}n_{\text{b}}}{\varepsilon_{0}m_{\text{b}}}$. In particular, the response to a static bias (${\bf{E}}={{\bf{E}}_{0}}$) is given by:
@@
$\mathbf{P}_{0}$ $=\varepsilon_{0}\sum_{j}\frac{\omega_{\text{b}}^{2}}{\omega_{0j}^{2}}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}\cdot\mathbf{E}_{0}.$ (22b)
@@
$\hskip 85.35826pt\varepsilon_{0}\omega_{\text{b}}^{2}\left({\mathbf{E}_{\omega}}+qn_{0}\sum_{jlr}v_{0r}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{l}\cdot\left[\left(a_{jlr}+a_{jrl}\right){\mathbf{v}_{\omega}}-i\omega b_{jrl}{\mathbf{P}_{\omega}}\right]\right)$
@@
$\mathbf{\overline{\varepsilon}}=\mathbf{\overline{\varepsilon}}_{\text{L}}+\mathbf{\overline{\chi}}_{\text{f-b}}+\mathbf{\overline{\chi}}_{\text{f-f}},$ (25)
where $\mathbf{\overline{\varepsilon}}_{\text{L}}$ determines the material response without the bias,
$\mathbf{\overline{\varepsilon}}_{\text{L}}$ $=\sum_{j}\varepsilon_{jj}^{0}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}$ (26a)
$\varepsilon_{jj}^{0}(\omega)$ $=1-\frac{\omega_{p}^{2}}{\omega\left(\omega+i\Gamma\right)}+\frac{\omega_{\text{b}}^{2}}{\omega_{0j}^{2}-\omega^{2}-i\omega\gamma}$ (26b)
and $\mathbf{\overline{\chi}}_{\text{f-b}}$ and $\mathbf{\overline{\chi}}_{\text{f-f}}$ represent the electro-optic response due to the interaction of bound and free electrons and due to the “anomalous velocity”, respectively:
$\mathbf{\overline{\chi}}_{\text{f-b}}$ $=-i\sum_{jlr}\left[g_{lr}a_{ljr}-g_{jr}\left(a_{jlr}+a_{jrl}\right)-f_{jlr}b_{jrl}\right]\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{l}$ (27a)
$\mathbf{\overline{\chi}}_{\text{f-f}}$ $=\frac{{{{\mathbf{\overline{\sigma}}}_{{\text{EO}}}}}}{{-i\omega{\varepsilon_{0}}}}.$ (27b)
@@
$f_{jlr}(\omega)$ $=-\varepsilon_{0}qn_{0}v_{0r}\omega\frac{\omega_{\text{b}}^{2}}{\omega_{0j}^{2}-\omega\left(\omega+i\gamma\right)}\frac{\omega_{\text{b}}^{2}}{\omega_{0l}^{2}-\omega\left(\omega+i\gamma\right)},$ (28a)
$g_{jr}(\omega)$ $=\varepsilon_{0}v_{0r}\frac{\omega_{p}^{2}\omega_{\text{b}}^{2}}{\left(\omega+i\Gamma\right)\left[\omega_{0j}^{2}-\omega\left(\omega+i\gamma\right)\right]}.$ (28b)
The linearized permittivity (25) is a dispersive function that in the absence of coupling ($a_{ijk}=b_{ijk}=0$ for all indices and $\mathbf{\overline{\zeta}}=0$) reduces to a diagonal matrix. As seen, the electro-optic effect may induce non-diagonal terms in the permittivity tensor. The term $\mathbf{\overline{\chi}}_{\text{f-b}}$ represents the part of the electro-optic response due to the interband effects that are not captured by the usual semiclassical calculation rooted on the Berry dipole ($\mathbf{\overline{\chi}}_{\text{f-f}}$). Note that $\mathbf{\overline{\chi}}_{\text{f-f}}$ has a pole in the static limit, whereas $\mathbf{\overline{\chi}}_{\text{f-b}}$ is finite in the same limit, confirming that the interband contribution is mostly relevant at high frequencies.
Remarkably, as discussed in the next section, for low-symmetry materials such non-diagonal terms may result in a permittivity tensor lacking transpose symmetry and may give rise to nonreciprocity and optical gain. In the rest of the article, we focus our analysis in the free and bound electrons (interband) interactions, and thereby we will ignore possible anomalous velocity contributions $\left(\mathbf{\overline{\zeta}}=0\right)$.